OpenStudy (anonymous):

Given the area of the left piston is 4m2, what is the area of the right piston (under the 600-kg mass)?

OpenStudy (anonymous):

that's the worksheet that we got

OpenStudy (anonymous):

\[P _{1}=P _{2} \] P=F/A F=Force, A=Area F=N=mg \[F _{1}/A _{1}=F _{2}/A _{2}\] \[20*9.8/4=600*9.8/A_{2}\] \[A _{2}=30 m ^{2}\]

OpenStudy (anonymous):

sorrry \[A _{2}=120 m ^{2}\]